OpenStudy (anonymous):

4(m+9)=-2(m-5)+2 m=?

jigglypuff314 (jigglypuff314):

Hello @D_spazz and Welcome to OpenStudy! :) to solve for m from 4(m+9) = -2(m-5) + 2 we should first distribute what we can :) that way we get 4m + 36 = -2m + 10 + 2 do you see how I got that so far? :)

OpenStudy (anonymous):

thank you yea

jigglypuff314 (jigglypuff314):

then we can add "like terms" because 10 + 2 = 12 so 4m + 36 = -2m + 10 + 2 becomes 4m + 36 = -2m + 12 then we have to try to get the numbers with "m"s to one side 4m + 36 = -2m + 12 +2m +2m ---------------------- 6m + 36 = 12

jigglypuff314 (jigglypuff314):

from there you can try subtracting 36 from both sides what would you get? :)

OpenStudy (anonymous):

-4 thank you

jigglypuff314 (jigglypuff314):

that is correct ^_^ good job!